# Data-Efficient Operator Learning via Unsupervised Pretraining and In-Context Learning

### v1

@@ 3.1.4 Model Architectures
We consider two popular architectures for fair comparisons with previous works. These are encoder-decoder architectures designed to reconstruct the original input given partial observations, where the encoder maps observed unlabeled PDE data to a latent space, and the decoder reconstructs the original conditions.
@@
Fourier Neural Operator (FNO) targets learning PDE data in the Fourier space. The original model backbone (encoder) employs Fourier transform and learns lower Fourier modes with linear transforms. The FNO backbone outputs features back to the spatial domain (i.e., the embeddings are on the pixel level). We refer readers to the original paper for details (Li et al., 2021a, b).
@@
Transformers, which mainly employ self-attention and linear transform blocks, have shown promise in both NLP and CV (Vaswani et al., 2017; Dosovitskiy et al., 2020). Different from FNO, which directly operates on grids, transformers tokenize and group grids into patches, i.e., each tokenized patch embeds a local neighborhood of subgrids. We follow the 3D transformer architecture of Video-MAE (Tong et al., 2022; McCabe et al., 2023). Our encoder embeds patches by a linear projection with added positional embeddings (just as in a standard ViT), and it then processes the resulting set via a series of Transformer blocks. For the transformer, the unlabeled PDE data are randomly masked at the patch level.
  • Pretraining: For the transformer encoder, we only apply it on the subset of tokens that are visible (i.e., unmasked patches), and masked patches are removed. This allows us to train encoders efficiently. The input to the MAE decoder is the full set of tokens consisting of (i) encoded visible patches, and (ii) the mask token. The mask token is a shared and learned vector that indicates the presence of a missing patch to be predicted. We add positional embeddings to all tokens in this full set. Without this, mask tokens would have no information about their location in the input. Following (He et al., 2022; Tong et al., 2022), we adopt an asymmetric design where the decoder is more lightweight (shallower and narrower) than the encoder.
@@ 3.2 In-Context Learning
Out-of-distribution (OOD) generalization is a critical technical challenge, not only in SciML but also across multiple domains in AI for science (Zhang et al., 2023). To improve the OOD generalizability of neural operators and to reduce the extra effort of downstream fine-tuning, the following inference paradigm has been proposed: given a query input, the model is also provided with a few supporting examples (dubbed “demos”), together with their ground-truth solutions, to make the final prediction. This approach enables the “open-set” generalization of the model to make predictions on unseen samples.
Originally, in the literature on few-shot learning (Zhang et al., 2018, 2020; Min et al., 2021; Shi et al., 2022; Kang and Cho, 2022; Liu et al., 2022; Peng et al., 2023), people developed delicate architectures to find a correspondence between the input and the supporting examples. The purpose of this extra architecture/training design is twofold: first, find similarities between the target input and supporting examples; and second, aggregate labels of supporting examples for the final predictions. Recent ICL works (Yang et al., 2023a, b) on learning PDE data also adopt this strategy, with transformers and cross-attention layers.
However, the ICL of LLMs enables a different strategy. The pretraining is still standard and simple (next/masked token prediction), without additional training costs. During inference, LLMs can auto-regressively take any number of few-shot examples, finding similarities between tokens in few-shot examples and those in the target query (via self-attention), and then generate responses by aggregating embeddings of tokens in few-shot examples. This ICL strategy used in LLM is highly scalable and training-efficient.
Motivated by this, we propose our simple ICL for neural operators. We describe our method in Algorithm 1.
@@
We find spatial-wise and temporal-wise similar demos by calculating their distance in the output space. That means, for two input locations over the spatial and temporal resolution, if we find their outputs of the trained neural operator similar, then we treat them as similar samples. We assume demos share the same distribution of physical parameters with the query.
@@
1 Data resolution: x-axis ($W$), y-axis ($H$), temporal steps ($T$), channel dimensions for the input ($C_{in}$) and solution ($C_{out}$).
2 Input: Query input ($x\in\mathds{R}^{H\times W\times C_{in}}$). Paired unlabeled PDE data ($X\in\mathds{R}^{J\times H\times W\times C_{in}}$) and solutions ($Y\in\mathds{R}^{J\times H\times W\times T\times C_{out}}$) as $J$ demos. Trained Neural Operator Model $\mathcal{M}$. TopK ($k$) demo solutions to aggregate.
3 $\hat{y}=\mathcal{M}(x)$triangleright Shape: $H\times W\times T\times C_{out}$
@@ 4 $\hat{Y}=\mathcal{M}(X)$triangleright Shape: $J\times H\times W\times T\times C_{out}$
5 $\hat{\delta}=\hat{y}\text{.reshape}(-1,1,C_{out})-\hat{Y}\text{.reshape}(1,-1,C_{out})$triangleright Shape: $H\times W\times T\times(J\cdot H\cdot W\cdot T)\times C_{out}$
@@
8 $\hat{y}_{icl}=$ take_along_dim($Y$.reshape$(-1,C_{out})$, index) $\triangleright$ Shape: $H\times W\times T\times C_{out}\times k$. Spatial and temporal aggregation of solutions from similar demos.
@@ 4 Empirical Results
To illustrate the benefits of our approach, we perform empirical evaluations and address the following questions. First, does our unsupervised pretraining (on unlabeled PDE data, followed by fine-tuning) outperform neural operators trained from scratch, while requiring fewer PDE simulations (Sec. 4.1)? Second, does our in-context learning help the model generalize better to OOD cases (Sec. 4.2)?
In our experiments, we trained models three times with different random seeds and provided error bars. For more experimental details, please refer to Appendix B. We also include ablation studies about pretraining hyperparameters in Appendix C. For visualizations of our pretraining, please refer to Appendix D.
Figure 3: Pretraining neural operators on unlabeled PDE data improves its performance and data efficiency on Poisson (a), Helmholtz (b), Reaction-Diffusion (c), and Navier-Stokes (d and e, with relative errors at different unrolled steps shown on f). “random init.”: models are trained from scratch with random initialization. “vision pretrained (SSv2)”: fine-tuning from the publicly available checkpoint for Video-MAE (pretrained on computer vision dataset SSV2 (Goyal et al., 2017) for video understanding). Savings of the number of simulated PDE data (when “random init.” achieves the best test error) are shown in red.
Figure 4: Benefits of unsupervised pretraining: 1) Universally reduced overfitting (i.e. smaller generalization gaps) on all PDEs (a to e); 2) Meaningful representations: pretraining Video-MAE with fixed encoder (red line) can extract meaningful features from unlabeled PDE data and outperform the baseline and the vision-pretrained model (f).
@@ 4.1 Unsupervised Pretraining Enables Data-Efficient Operator Learning
We first demonstrate that, by leveraging unsupervised pretraining, neural operators can achieve improved relative errors with less simulated data. See Figure 3. We can see that, compared with directly training neural operators from scratch (“random init.”), pretraining on unlabeled PDE data can help neural operators achieve better performance, which can further help reduce the amount of simulated data. Specifically, in our experiments, when we target achieving the best test error of the baseline (“random init.”), our method can save $5\times 10^{2}\sim 8\times 10^{5}$ simulated solutions across diverse PDE systems.
Among all PDEs, we find that Helmholtz (Figure 3 (b)) is the most challenging. Two curves failed to improve the error until we increased the number of simulated data points to over 1024. Meanwhile, the generalization gaps remain high (Figure 4 (b)), indicating low training errors. We suspect that learning on the Helmholtz equation may be exhibiting the grokking issue (Power et al., 2022; Zhu et al., 2024), where the network quickly memorizes the training data, but the improvement in generalizability is delayed.
@@
Pretraining on unlabeled PDE data is not the only way to save simulation costs. As pretraining is widely adopted in CV, pretrained checkpoints on vision data become publicly available. We choose to compare with Video-MAE (Tong et al., 2022) for state-of-the-art video understanding pretrained on CV dataset SSV2 (Goyal et al., 2017). As shown in Figure 3 (e), vision-pretrained Video-MAE can only outperform the random initialization with high volumes of simulated data, while its performance suffers when fine-tuned with limited simulations. In contrast, our unsupervised pretraining on unlabeled PDE data can save a significant amount of simulated data.
@@
Pretraining on unlabeled PDE data is beneficial beyond achieving better performance with less simulated PDE data. First, we find that pretraining on unlabeled PDE data can strongly regularize the model against overfitting, without the need for any early-stopping heuristics (our smaller generalization gaps in Figure 4 (a-e)). Second, unsupervised pretraining can also help models extract useful representations for subsequent operator learning. For example, in Figure 4 (f), we find that with even pre-extracted features (i.e., model’s encoder is fixed during fine-tuning), our neural operators can still outperform baselines (trained from scratch).
@@ 4.2 ICL Enables Data-Efficient OOD Generalization
We now move to OOD settings, where models will be tested on PDE data simulated with physical parameters unseen during fine-tuning/training. Neural operators suffer from poor OOD generalization (Subramanian et al., 2023; Yang et al., 2023a). Traditionally, improvements heavily depend on further fine-tuning on simulated data, which requires extra simulation and training costs. We study the benefits of our test-time in-context learning. As shown in Figure 5, when we flexibly scale up the number of demos, we can keep improving FNO’s OOD generalization on diverse PDEs. Notably, we introduce zero training overhead: we keep the standard training pipeline, and our ICL can be seamlessly plugged in during OOD inference.
We show visualizations of in-context learning in Figure 5. In this visualization, we find that the range of numerical solutions (e.g., values in colorbars) predicted by in-context learning become closer to the target. Please refer to Appendix F for further discussions about the benefits of ICL.
@@ 5 Conclusion
In this work, we focus on enhancing the data efficiency of solving PDEs using deep neural networks, particularly through unsupervised pretraining and ICL methods. Current machine learning methods for PDEs are data-intensive, which can be expensive to obtain. Our key contributions include the introduction of unsupervised pretraining for neural operator learning and a flexible ICL approach that enhances OOD generalization without incurring extra training costs. Through extensive empirical evaluations, we demonstrate that our method is not only data-efficient but also more generalizable compared to existing approaches. We anticipate that our work would inspire the SciML community to explore more on solving the heavy simulation costs and unsatisfactory OOD generalization of neural operators.

### v2

@@ 3.1.4 Model Architectures
We consider two popular architectures for fair comparisons with previous works. These are encoder-decoder architectures designed to reconstruct the original input given partial observations, where the encoder maps observed unlabeled PDE data to a latent space, and the decoder reconstructs the original conditions. We include visualizations of these architectures in Appendix D.
@@
Fourier Neural Operator (FNO) targets learning PDE data in the Fourier space. The original model backbone (encoder) employs Fourier transform and learns lower Fourier modes with linear transforms. The FNO backbone outputs features back to the spatial domain (i.e., the embeddings are on the pixel level). We refer readers to the original paper for details [4, 9].
@@
Transformers, which mainly employ self-attention and linear transform blocks, have shown promise in both NLP and CV [63, 64]. Different from FNO, which directly operates on grids, transformers tokenize and group grids into patches, i.e., each tokenized patch embeds a local neighborhood of subgrids. We follow the 3D transformer architecture of Video-MAE [65, 47]. Our encoder embeds patches by a linear projection with added positional embeddings (just as in a standard ViT), and it then processes the resulting set via a series of Transformer blocks. For the transformer, the unlabeled PDE data are randomly masked at the patch level.
  • Pretraining: For the transformer encoder, we only apply it on the subset of tokens that are visible (i.e., unmasked patches), and masked patches are removed. This allows us to train encoders efficiently. The input to the MAE decoder is the full set of tokens consisting of (i) encoded visible patches, and (ii) the mask token. The mask token is a shared and learned vector that indicates the presence of a missing patch to be predicted. We add positional embeddings to all tokens in this full set. Without this, mask tokens would have no information about their location in the input. Following [19, 65], we adopt an asymmetric design where the decoder is more lightweight (shallower and narrower) than the encoder.
@@ 3.2 In-Context Learning
Out-of-distribution (OOD) generalization is a critical technical challenge, not only in SciML but also across multiple domains in AI for science [38]. To improve the OOD generalizability of neural operators and to reduce the extra effort of downstream fine-tuning, the following inference paradigm has been proposed: given a query input, the model is also provided with a few supporting examples (dubbed “demos”), together with their ground-truth solutions, to make the final prediction. This approach enables the “open-set” generalization of the model to make predictions on unseen samples.
Originally, in the literature on few-shot learning [66, 49, 67, 68, 69, 70, 71], people developed delicate architectures to find a correspondence between the input and the supporting examples. The purpose of this extra architecture/training design is twofold: first, find similarities between the target input and supporting examples; and second, aggregate labels of supporting examples for the final predictions. Recent ICL works [24, 25, 26] on learning PDE data also adopt this strategy, with transformers and cross-attention layers.
However, the ICL of LLMs (large language models) enables a different strategy. The pretraining is still standard and simple (next/masked token prediction), without additional training costs. During inference, LLMs can auto-regressively take any number of few-shot examples, finding similarities between tokens in few-shot examples and those in the target query (via self-attention), and then generate responses by aggregating embeddings of tokens in few-shot examples. This ICL strategy used in LLM is highly scalable and training-efficient.
Motivated by this, we propose our simple ICL for neural operators with two steps (Algorithm 1).
@@
We find spatial-wise and temporal-wise similar demos by calculating their distance in the output space. That means, for two input locations over the spatial and temporal domains, if we find their outputs of the trained neural operator similar, then we treat them as similar samples. Following [24, 25], we assume demos share the same distribution of physical parameters with the query.
@@
1
2Data resolution: x-axis ($W$), y-axis ($H$), output temporal steps ($T$), output channel dimensions for the solution ($C_{out}$).
3 Input: Query input ($x$). Paired unlabeled PDE data ($X$) and solutions ($Y\in\mathds{R}^{J\times H\times W\times T\times C_{out}}$) as $J$ demos. Trained Neural Operator Model $\mathcal{M}$. TopK ($k$) demo solutions to aggregate.
4
[…]
@@ 6 $\hat{Y}=\mathcal{M}(X)$triangleright Shape: $J\times H\times W\times T\times C_{out}$
7 $\hat{\delta}=\hat{y}\text{.reshape}(-1,1,C_{out})-\hat{Y}\text{.reshape}(1,-1,C_{out})$triangleright Query-Demo Distance. Shape: $H\times W\times T\times(J\cdot H\cdot W\cdot T)\times C_{out}$
@@
10 $\hat{y}_{icl}=$ take_along_dim($Y$.reshape$(-1,C_{out})$, index) $\triangleright$ Spatial-wise and temporal-wise aggregation of solutions from similar demos. Shape: $H\times W\times T\times C_{out}\times k$.
@@ 4 Empirical Results
To illustrate the benefits of our approach, we perform empirical evaluations on both PDE benchmarks and real-world observations. Most importantly, our unsupervised pretraining (on unlabeled PDE data, followed by fine-tuning) outperforms neural operators trained from scratch, while requiring fewer PDE simulations (Sec. 4.1 and Sec. 4.2). Moreover, our in-context learning can help the model generalize better to OOD cases (Sec. 4.3). In our experiments, we trained models three times with different random seeds for statistical significance. For more experimental details, please refer to Appendix B. We also include ablation studies about pretraining hyperparameters in Appendix G. For visualizations of our pretraining, please refer to Appendix H.
Figure 3: Pretraining neural operators on unlabeled PDE data improves its performance and data efficiency on Poisson (a), Helmholtz (b), Reaction-Diffusion (c), and Navier-Stokes (d and e, with relative errors at different unrolled steps shown on f). “random init.”: models are trained from scratch with random initialization. “vision pretrained (SSv2)”: fine-tuning from the publicly available checkpoint for Video-MAE (pretrained on computer vision dataset SSV2 [72] for video understanding). Savings of the number of simulated PDE data (when “random init.” achieves the best test error) are shown in red.
@@ 4.1 Unsupervised Pretraining Enables Data-Efficient Operator Learning
We first demonstrate that, by leveraging unsupervised pretraining, neural operators can achieve improved errors with less simulated data. In Figure 3, compared with directly training neural operators from scratch (“random init.”), pretraining on unlabeled PDE data can help neural operators achieve better performance, which can further help reduce the amount of simulated data. Specifically, in our experiments, when we target achieving the best test error of the baseline (“random init.”), our method can save $5\times 10^{2}\sim 8\times 10^{5}$ simulated solutions across diverse PDE systems.
Among all PDEs, we find that Helmholtz (Figure 3 (b)) is the most challenging. Both two curves failed to improve the error until we increased the number of simulated data points to over 1024. Meanwhile, the generalization gaps remain high (Figure 9 (b)), indicating low training errors. We suspect that learning on the Helmholtz equation may be exhibiting the grokking issue [73, 74], where the network quickly memorizes the training data, but the improvement in generalizability is delayed.
@@
Pretraining on unlabeled PDE data is not the only way to save simulation costs. As pretraining is widely adopted in CV, pretrained checkpoints on vision data become publicly available and are ready for fine-tuning. We choose to compare with Video-MAE [65] for state-of-the-art video understanding pretrained on SSV2 [72]. As shown in Figure 3 (e), vision-pretrained Video-MAE can only outperform the random initialization with high volumes of simulated data, while its performance suffers when fine-tuned with limited simulations. In contrast, our unsupervised pretraining on unlabeled PDE data can save a significant amount of simulated data.
@@
Pretraining on unlabeled PDE data is beneficial beyond achieving better performance with reduced simulations. First, when training on extremely low volumes of data, neural operators tend to overfit, resulting in poor generalization. In Figure 4 left, pretraining on unlabeled PDE data can reduce the generalization gap (testing error -- training error). In Figure 9, we further show that better generalization gaps persist across all PDEs we studied. Second, pretraining on unlabeled PDE data can lead to faster convergence during fine-tuning. Unlike standard random initializations from Gaussian distributions, pretraining on unlabeled PDE data will provide neural operators with domain-adapted initializations and facilitate a much faster convergence rate, as shown in Figure 4 middle [^2] Training curves when the number of training samples of 14760.. Third, unsupervised pretraining can also help models extract useful representations for subsequent operator learning. From Figure 4 on the right, we find that even with pre-extracted features (i.e., fixed encoder and only fine-tuned decoder), our neural operators can still outperform the baselines where both the encoder and decoder are updated during fine-tuning.
Figure 4: Benefits of our unsupervised pretraining. Reduced overfitting (left): our method consistently leads to smaller generalization gaps (test error -- training error) across all PDEs we studied (Fig. 9). Faster convergence (middle): our unsupervised pretraining can accelerate model convergence than both random initialization and vision-pretrained checkpoint. Meaningful representations (right): fine-tuning Video-MAE with fixed encoder (pretrained on unlabeled PDE data, red line) can extract meaningful features and outperform the baseline and the vision-pretrained model (both encoder and decoder are updated during fine-tuning).
Figure 5: For real-world scientific problems, pretraining neural operators on unlabeled PDE data improves its performance and data efficiency. Top (a-j) (visualizations): We show snapshot examples from ERA5 temperature (a, b) and ScalarFlow (c, d) at different temporal steps; and also an example of airfoil mask, velocities, and pressure (e-j[75]. Bottom (test errors): We study VideoMAE [65] pretrained with unlabeled snapshots (no temporal dynamics), and then fine-tune across different numbers of temporal snapshots on ERA5 (bottom left) and ScalarFlow (bottom middle). We also pretrain 2D-FNO [4] on freestream velocities and fine-tune on time-independent steady-state airflow pressure and velocities (bottom right). “random init.”: models are trained from scratch with random initialization.
4.2 More Comprehensive Experiments on Real-World Data
In this section, we move to a broader range of benchmarks. We will study real-world and noisy data instead of toy datasets, providing even more comprehensive experiments. These benchmarks are widely studied in previous works [76, 77, 47].
Model and Training.
For time-dependent ERA5 (Sec. 4.2.1) and ScalarFlow (Sec. 4.2.2) datasets, we adopt the same VideoMAE architecture [65] used in Sec. 4.1. We use 15 consecutive temporal snapshots to forecast the next time step. We train VideoMAE with Adam, with other hyperparameters the same as in Table 2 column “N.S. (PDEBench)”. For the time-independent steady-state Airfoil dataset (Sec. 4.2.3), we adopt the 2D-FNO architecture. We train 2D-FNO with Adam, with other hyperparameters the same as in Table 2 column “Poisson”.
4.2.1 ECMWF Reanalysis v5 (ERA5)
The ECMWF (European Centre for Medium-Range Weather Forecasts) offers ERA5, a public extensive dataset [78], which delivers hourly data on multiple atmospheric variables spanning from 1979 to today. ERA5 represents a type of atmospheric reanalysis dataset [79], integrating observations from a variety of measurement sources with numerical models through data assimilation [80]. Essentially, it reconstructs the most accurate estimation of the Earth’s atmospheric conditions over time. This dataset has been extensively utilized in prior scientific machine learning (SciML) studies [76, 59].
In this experiment, we focus on forecasting the important and challenging temperature atmospheric variable. We utilize data from 2006 to 2015, with the snapshots taken every 6 hours and a spatial resolution of $360\times 360$. The total number of snapshots is 14600. We apply the mean-standard deviation normalization to the data and downsample the snapshots to a spatial resolution of $180\times 180$. We split 75% of the data for pretrain and 25% for finetune. For each split, we further separate 80% of the data for training, 10% for validation, and 10% for testing.
Results.
We show results in Figure 5 (bottom left). Compared with directly training neural operators from scratch (“random init.”), pretraining on 2D snapshots without temporal dynamics can help neural operators achieve better performance and reduce the amount of simulated data.
4.2.2 ScalarFlow
ScalarFlow [81] is a dataset of reconstructions of real-world smoke plumes. It assembles the first large-scale dataset of realistic turbulent flows. The availability of a large, volumetric data set opens up a wide range of applications, including re-simulations, novel visualization, and metric evaluations. The dataset contains 104 real-world smoke flow density reconstructions. Each reconstruction is captured from five different viewpoints for 150 temporal frames spanning 2.5 seconds. The original spatial resolution is $1062\times 600$. We crop the snapshots to $900\times 600$ to remove the padding and background. We remove the first 15 timeframes for each simulation to avoid the initial transient phase at the beginning of the smoke flow generation. We then downscale the snapshots to $180\times 120$ and apply mean-standard deviation normalization to the data. With a total of 70200 snapshots, we split 80% of the data for pretrain and 20% for finetune. For each split, we further separate 80% of the data for training, 10% for validation, and 10% for testing.
Results.
We show our results in Figure 5 (bottom middle). Pretraining neural operators on unlabeled PDE data, specifically 2D snapshots that lack temporal dynamics, as opposed to training them from the beginning with no prior knowledge (“random init.”), can enhance their performance even on smoke flow data with strong turbulence.
4.2.3 Airfoil
Airfoil data introduced in [75] is a large-scale dataset that contains the pressure and velocity simulations of the flow around the airfoils. This dataset has 53880 samples for training and 90 samples for testing. We split 80% of the training data for pretrain and 20% for finetune. Each sample contains 6 channels. The three input channels include the binary spatial mask of the airfoil and the initial velocity components of the freestream condition in the x and y directions. The output channels include the pressure and the x and y velocity components from the simulation at the steady state. The original spatial resolution of each sample is $128\times 128$. We crop the snapshots to $96\times 45$ to remove the background. We then apply min-max normalization channel-wise to the sample.
Results.
We show our results in Figure 5 (bottom right). For this time-independent scenario, pretraining on unlabeled data (i.e., the freestream velocities) is still beneficial for FNO to achieve more accurate predictions of the steady-state pressure and velocity flow around airfoils.
@@ 4.3 ICL Enables Data-Efficient OOD Generalization
We now move to OOD settings, where models will be tested on PDE data simulated with physical parameters unseen during fine-tuning/training. We include how to simulate OOD samples in Appendix B.2. Neural operators suffer from poor OOD generalization [11, 24]. Traditionally, improvements heavily depend on further fine-tuning on simulated data, which requires extra simulation and training costs. We study the benefits of our test-time in-context learning. As shown in Figure 6, when we flexibly scale up the number of demos, we can keep improving FNO’s OOD generalization on diverse PDEs. We follow [24, 25] that demos are randomly sampled from the same distribution used to generate the OOD test set. When the number of demos is 0, we have the baseline in the OOD setting. Notably, we introduce zero training overhead: we keep the standard training pipeline, and our ICL can be seamlessly plugged in during OOD inference.
We show visualizations of in-context learning in Figure 6. In this visualization, we find that the range of numerical solutions (e.g., values in colorbars) predicted by in-context learning becomes closer to the target. Please refer to Appendix I for further discussions about the benefits of ICL.
@@ 5 Conclusion
In this work, we focus on enhancing the data efficiency of solving PDEs using deep networks, particularly through unsupervised pretraining and ICL methods. Our key contributions include the introduction of unsupervised pretraining for operator learning and a flexible ICL approach that enhances OOD generalization without incurring extra training costs. Through extensive evaluations, we demonstrate that our method is not only data-efficient but also more generalizable compared to existing approaches. We anticipate that our work would inspire the SciML community to explore more on solving the heavy simulation costs and unsatisfactory OOD generalization of neural operators. Current limitations include: 1) more proxy tasks could be further designed; 2) larger SciML foundation models could be studied; 3) more physical constraints (e.g. the conservation law) could be considered. These limitations are beyond the scope of this work, and we expect to address them in future research.